OpenStudy (anonymous):

write an equation in slope-intercept form of the line that is perpendicular to the graph of each equation and passes through the given point. 1. y=-5x+1; (2,-1) 2. 3x-4y=2; (6,0)

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

compare this with y=mx+c your line would be y = -(1/m)x + b put those values of x and y and get value of b.

OpenStudy (anonymous):

yea i know that part its just some part that am stuck at cause of the fraction

OpenStudy (experimentx):

y=-5x+1; y=mx+c -------- m=-5 ---------------------- your line would be y = -(1/(-5))x + b y = (1/5)x + b put this (x,y) = (2,-1) ------------ (-1) = (1/5)(2) + b b = -1 - 2/5 = -7/5 so you line is y = x/5 -7/5

OpenStudy (anonymous):

yea if it wasn't in the slope of a perpendicular, because the lines are negative reciprocal of each other

OpenStudy (anonymous):

so for example y=-7x+2 the new perpendicular slope would be 1/7

OpenStudy (experimentx):

yeah ...just change it 3x-4y=2 4y = 3x - 2 y = 3/4 x - 2/4 y = mx + c m = 3/4 and c=-2/4

OpenStudy (experimentx):

yes ... it would be

OpenStudy (anonymous):

wait is that the working out for the second one and do understand what am saying cause i might not explain things so good

OpenStudy (experimentx):

yeah ... I"m understading.

OpenStudy (anonymous):

ok, so are you gonna work it out step by step for me?

OpenStudy (experimentx):

this is similar to the first one. try this one ... for pratice. I'll see if you are correct or not.

OpenStudy (anonymous):

ok well i got a different answer i got -4y=-3x+2 i subtracted 3x from both side

OpenStudy (experimentx):

you should get y = 8-(4/3)x

OpenStudy (anonymous):

ok i'll try and see how to do that, even though am not sure how to

OpenStudy (experimentx):

3x-4y=2 4y = 3x - 2 y = 3/4 x - 2/4 y = mx + c m = 3/4 and c=-2/4 so -1/m = -4/3 so your line should be y = -4/3 x + b ---------------- put (x,y) = (6,0) and get the value of b. put that value of b in above equation and get your answer.

OpenStudy (anonymous):

ok thanks, and that's the same way for the first one too